FEL Upgrade Project Weekly Brief July 29 - August 2, 2002

 

 

Highlights:

Our first dipole magnet for the FEL Upgrade, one of the new injector magnets (GV), is expected to be delivered to us by air freight today. This magnet will be the first new dipole to go through the magnetic measurement for qualification and then onto installation in the injector area.

The I&C team has successfully demonstrated a new video capture system that will allow us to automate electron beam brightness (i.e., emittance) measurements in the FEL Upgrade driver accelerator.

Management:

We are pleased to recognize the arrival of $1.517M of funds from the Air Force Research Lab to fund the second phase of activities for the UV FEL Upgrade and the collaboration with the Aerospace Corporation for development of a UV laser processing tool. We thank John Eric from AFRL and Wayne Skinner from DOE for their assistance in processing the documentation for the funds.

G. Neil and F. Dylla attended a review of injector technologies for high power FEL development sponsored by the JTO Technical Working Group (TAWG) on FELs this week held at the facilities of AES, Inc. in Medford, NY. We presented a summary of experience with the IR Demo photogun which operated at 5mA, our improvements incorporated in the 10 mA FEL Upgrade gun, and a summary of our collaboration with AES on the development of a 100 mA gun for the proposed 100 kW FEL Upgrade.

Planning continued on finalizing labor and budget allocations for FY03 activities based on our best estimates of available FY03 funding for FEL tasks.

WBS 4 (Injector):

The 6-way cross and shield operator bellows were welded. The drawing package for the resistor is in review. Test fit up with a metal mirror is being done on the light box. The support tube for the ball cathode was mounted this week in the W&M ion implantation system. The first two of six implant runs were completed on Thursday. The second run had to be performed at 5 kV instead of 15 kV due to instabilities at the end of the current pulse, caused probably by the misalignment of the tube in the chamber's sidearm. The mount has been repositioned to continue with the rest of the implant runs next week.

Gun HVPS - The connector flanges, a frame for the electronics, relief valve, SF6 piping, and a lighter weight castor are needed for the tank. The assembly of the HVPS will start next week. Additional material is on order for the 2nd conditioning resistor. The 1st conditioning resistor was repaired this week and can be used as a model and a spare. It measures 186 M Ohms at low voltage.

WBS 6 (RF):

Zone 4 - Requisitions were written and signed today to purchase the flexible waveguides necessary to complete the installation to the second cryomodule. Several additional sections of waveguide will also need to be procured.

Quarter HVPS - The contactor, heater code, and HV code was successfully tested this week. The 480 VAC will be connected next week and the unit will be readied for HV testing. The software effort to operate remotely has started.

Quarter Klystrons - Still need some minor wiring before HV testing can begin. This is also planned for next week.

WBS 8 (Instrumentation):

A great deal of progress has been made on DC power this week. There was a design review for the 100 Amp sextapole and octapole power supplies. The design will be based on a commercial switcher and an external current transducer with 5 ppm stability. An interface chassis will be built that contains a switch (for bipolar) and a closed loop regulator. The (deeper) racks for these supplies arrived, were installed, and had the three phase 208 plumbed in. The drawings are being finalized for the higher current cable runs. The 100Amp cable is on order and the 535MCM cable will be reused. The wiring harness's for the DB/DJ correctors have been fabricated and will be installed on the girders as they arrive in the vault. Termination of the magnet trim cables in junction boxes 4 and 5 has been completed, rung out and labeled. The conversion of the trim racks from CAMAC based controls to a serial interface is in progress.

Both the Switched Electrode Electronics (SEE) and the four channel Beam Position Monitor (BPM) cables were reinstalled this week. There are an additional 8 BPMs so more cable will be ordered. The RF modules for the SEE's have been turned over to the I&C group in the main machine for testing and calibration.

Drive Laser Pulse Controller (DLPC) - A second timer card (F0080) has been implemented into iocfel10 for the drive laser pulse controller upgrade. It's function is to provide independent range and resolution control of the macropulse width which is therefore independent of the macropulse rep-rate. The card pair has been tested to recreate all the old modes of operation (synchronous and asynchronous to 60Hz) and works well. The focus now is to integrate the CO305 card and MPS functions.

Beam Loss Monitor (BLMs) - The BLM front card (F0151) PCB layout has been completed and is ready for fabrication (later today). The majority of the parts are in and there are no long lead times for any of the remaining items.

 

 

Effort continued on the HV interlock chassis. The design is complete and awaiting EECAD for

schematic and then PCB layout. The new design will combine functions from three different chassis from the IRDEMO

Progress also continues on learning how to efficiently use the Motorola 807 DSP. The application that is being worked is processing the signals from a 128 segment line scan camera and outputting the FEL spectrum (using a grating). We had started with few ADC programming examples available from Motorola, to understand the working of on-board ADCs. There are 2 options. While talking to the board, firstly generating code using Motorola provided drivers and libraries (called the SDK tools). Secondly generating code without using Motorola drivers or libraries. (1)The previous modules (presently running them as separate programs) developed for Clkin_gen and Beamsync_Pulse(o/p) have now been combined and run using SDK. (2)Rewritting the programs separately without using the SDK tools now. This would ensure that the execution would be faster and code more readable by the user if changes are required.

Lastly the Video acquisition system is up and running under Linux. The software is being worked on to create the tools for an automated emittance measurement system. We will have the PC running Linux acting as an IOC with no VXWorks license fees. The automatic calibration (magnification) and background subtraction routine are being coded now.

WBS 9 (Transport):

Dipoles

Optical Chicane Dipoles (GW)

At Magnet Enterprises International (MEI) in Oakland CA. the next three GWs were shipped in the middle the week.

The remaining 6 cores of the second batch of 6 Magnets for UV are ready for gap checking. The instrument will arrive next week.

All coils are potted.

Injector Dipoles (DU/DV)

Wang NMR glued the shim materials to the some GU cores.

The GVs were assembled ad the first was shipped by airfreight.

The missing GV field clamp parts were found at Master Machine, sent to Wang and became part of the assembly.

Arc 180 Degree Dipoles (GY)

At Wang NMR, the fourth GY Coil is nearing completion of winding.

GY Coil Potting fixture remains in fabrication.

We are resolving the contractual problems resulting from having to wind a fifth coil.

Arc Bend, Reverse Bend Dipoles (GQ, GX)

Process Equipment Co. (PECo) assembled the remaining three GQ cores and will ship early next week. The GX cores are on track for delivery two weeks later.

At Wang NMR, potting fixtures remain in preparation for the GQ and GX coils.

Quadrupoles

Trim Quad (QT)

The first unit passed temperature test with added cooling of the side coils. We generated an engineering change order with the details so that Milhous can add this feature to the existing units.

Sextupole (SF)

DULY research will finish the Sextupoles over the next two weeks.

Octupole (OT)

DULY started design of the Octupole.

Beam Line and Vacuum

Master Machine started machining the flanges of the Arc Vacuum Chambers and received the material. They are working with a vendor with a 5 axis water jet cutting machine to cut out the parts. We inspected a sample piece of stainless steel that had been water jet cut and determined that there was minimal contamination from the cut method and that in addition, the style of the weld joint would lead to minimal contamination of the vacuum surfaces. We concluded that water jet cutting for these chambers is acceptable.

The Jlab Shop started tack welding the X Chamber together and received material to start fabrication of the four chambers for the optical chicane region.

Design of the regions around the wiggler is 70 % complete.

The stands for sextupoles and trim quads are in procurement.

Design of the Octupole stand started.

The first meeting was held to start design of the recirculation dump region.

WBS 11 (Optics):

This week's work on the optical cavity involved replacing an uncoated leadscrew/ Si:bronze nut combination with a leadscrew that was sputter coated with silver. This combination had greatly reduced torque, but revealed a sensitivity to a subtle bowing of the leadscrew. We have a solution that resolves the problem and it is being implemented. We have decided to use the current rotary feedthroughs (c.f.last week's report) while discussions are underway with Lab staff who have prior experience in this area to quickly find another vendor. We received modified flexural pivots for the yaw plate, and it now moves smoothly.

We met with a representative of the LVDT vendor, and discussed purchasing units that can be baked to a higher temperature (yielding lower outgassing rates. The rest of the linear actuators for the cavity, and most of the optical transport has been placed. We held a meeting with I&C staff to determine the division of labor and responsibility for wiring of the optical cavity. We have the new position sensitive detectors for the OCMMS and are beginning to prototype the electronics.

Our tests of the prototype backplane-cooled mirror assembly in the mirror test stand showed that the design does a good job of dissipating thermal loading higher than anticipated. The total wavefront aberration at the highest loading we tried, 65 W/sqr. cm. was ~ 1 wave at 633 nm. The aberration is linear with the absorbed power, and at the specified power output and loading, is only ~ 0.2 waves. Since we wish to make the optical transport readily upgradeable to higher FEL outputs, this suggests we make the substrate thinner, and we will discuss this with various vendors to see if that is feasible.

We continue to work to resolve the power instability that is plaguing the ultrafast laser system. If nothing else, one simply has to leave it on for a few hours, but it should perform better. Work on the diagnostics suite continues. We received a BBO crystal to generate UV for laser induced damage testing. This should begin today. An order was placed for the rest of the new drive laser transport mirrors.